# Robust Blockwise Random Pivoting:

### v3

@@ 3.1 Random Pivoting
For any $\epsilon>0$ and $r\in[n]$, the skeleton subset $S$ selected by Algorithm 1 provides $\mathbb{E}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leq\left(1+\epsilon\right)\left\|\mathbf{X}-{\mathbf{X}}_{\langle r\rangle}\right\|_{F}^{2}$ when333 Without strictly specifying, we generally assume that $\epsilon>0$ is small enough such that the $(r,\epsilon)$-ID is a non-trivial approximation, i.e. $\epsilon<\min\left\{1/\eta_{r},2^{r}\right\}$ such that both logarithmic terms are positive. Otherwise, the skeleton complexity guarantee still holds by replacing any negative logarithmic terms with zeros.
$k=\left|S\right|\geq\frac{r}{\epsilon}+r\cdot\min\left\{\log\left(\frac{1}{\epsilon\eta_{r}}\right),1+\log\left(\frac{2^{r}}{\epsilon}\right)\right\}.$
From Proposition 3.3, we notice that sequential random pivoting almost matches the nearly optimal skeleton complexity guarantee of DPP (Equation 7) up to some logarithmic factors $\log\left(\frac{1}{\epsilon\eta_{r}}\right)$ on $r$. Further, the appealing skeleton complexity of random pivoting is empirically demonstrated in the form of RPCholesky for column Nyström approximation in [5].
@@
Although random pivoting leverages the strength of both adaptiveness and randomness and achieves an attractive skeleton complexity with affordable asymptotic complexity, the lack of parallelizability due to its sequential nature becomes a critical concern when considering empirical runtime.
@@
Given a sequence of matrix operations, it is well-known that an appropriate implementation using Level-3 BLAS, or matrix-matrix, operations will run more efficiently on modern processors than an optimal implementation using Level-2 or Level-1 BLAS [3]. This is largely due to the greater potential for the Level-3 BLAS to make more efficient use of memory caching in the processor. Notice that the computational bottleneck of Algorithm 1 with asymptotic complexity $O\left(ndk\right)$ (or $O\left(ndk\right)$) consists of $k$ matrix-vector multiplications (Level-2 BLAS) with $\mathbf{X}$. This leads to the considerable slowdown of Algorithm 1 in practice.
@@ 3.2 Sketchy Pivoting
Alternative to random pivoting, randomness can be combined with adaptiveness through randomized linear embedding (also known as sketching), which leads to the sketchy pivoting methods [34, 48, 16, 19]. As given in Algorithm 3, the general framework of sketchy pivoting [16, Algorithm 1] consists of two stages.
Algorithm 3 Sketchy pivoting [48, 16] (SkLUPP/SkCPQR)
1:(a) $\mathbf{X}=\left[\mathbf{x}_{1},\cdots,\mathbf{x}_{n}\right]^{\top}\in\mathbb{R}^{n\times d}$, (b) Target rank $k\in\mathbb{N}$, (c) Sample size $l\geq k$, (d) Pivoting strategy: LUPP/CPQR
3:Draw a randomized linear embedding $\bm{\Omega}\in\mathbb{R}^{d\times l}$
4:$\mathbf{Y}\leftarrow\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$triangleright $O\left(ndk\right)$
5:if LUPP then
[…]
7:else if CPQR then
[…]
In the first stage, randomness is incorporated through sketching, which serves two purposes.
  • From the computational efficiency perspective, sketching reduces the data dimension from $d$ to $l=O(k)$444 For applications like low-rank approximations and skeleton selection, a small constant oversampling like $k+5\leq l\leq k+10$ is usually sufficient in practice. , while transferring the $O(ndk)$ computational bottleneck from the sequential greedy pivoting to the (input-sparsity-time and) parallelizable matrix-matrix multiplications.
@@
Common choices of randomized linear embeddings in the sketching stage include Gaussian embeddings [28] (with asymptotic complexity $O\left(ndk\right)$), subsampled randomized trigonometric transforms [50, 43, 46, 4] (with asymptotic complexity $O\left(nd\log(k)\right)$), and sparse embeddings [36, 37, 7, 47] (with asymptotic complexity $O\left(nd\right)$). We refer the interested readers to [26, 49, 35] for a general review of sketching.
In the second stage, adaptiveness is utilized as greedy pivoting on the sketched sample matrix $\mathbf{Y}=\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$ to select skeletons. Both LUPP [22, Section 3.4] and CPQR [22, Section 5.4.2] on $\mathbf{Y}$ cost $O(nk^{2})$ asymptotically, whereas LUPP is remarkably faster in practice due to its superior parallelizability [16]. Intuitively, applying greedy pivoting on the top of sketching can be viewed as an analog of the squared-norm sampling with a slightly different squared-norm-based distribution.
For example, for a Gaussian random matrix $\bm{\Omega}$ with i.i.d. entries from $\mathcal{N}\left(0,1/l\right)$, each row $\mathbf{y}_{i}=\bm{\Omega}^{\top}\mathbf{x}_{i}$ of $\mathbf{Y}$ is a Gaussian random vector with i.i.d. entries from $\mathcal{N}\left(\textbf{0},\left\|\mathbf{x}_{i}\right\|_{2}^{2}\mathbf{I}_{l}\right)$555 It is worth highlighting that although rows of $\mathbf{Y}$ consist of i.i.d. entries, entries in each column of $\mathbf{Y}$ are dependent. In particular, for all $j\in[l]$, $\mathrm{Cov}\left(\mathbf{Y}\left(:,j\right)\right)=\mathbf{X}\mathbf{D}^{-1}\mathbf{X}^{\top}$ where $\mathbf{D}=\mathop{\mathrm{diag}}\left(\left\|\mathbf{x}_{1}\right\|_{2}^{2},\cdots,\left\|\mathbf{x}_{n}\right\|_{2}^{2}\right)$. . When applying greedy squared-norm pivoting (CPQR) in the second stage, the first pivot is selected according to $s_{1}\leftarrow\mathop{\mathrm{argmax}}_{i\in[n]}\left\|\mathbf{x}_{i}\right\|_{2}^{2}Z_{i}$ where every $Z_{i}=\sum_{j=1}^{l}Y_{ij}^{2}$ is a $\chi^{2}_{l}$ random variable but $\left\{Z_{i}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ i\in[n]\right\}$ are dependent. The intuition behind sketchy pivoting is similar to random pivoting, but the randomness arises from the random variables $Z_{i}$, instead of sampling.
In the extreme scenarios, when the variance of $\left\{Z_{i}\right\}_{i\in[n]}$ is negligible in comparison to the squared norms $\left\{\left\|\mathbf{x}_{i}\right\|_{2}^{2}\right\}_{i\in[n]}$, sketchy pivoting behaves similarly to greedy pivoting. Meanwhile, when the variance of $\left\{Z_{i}\right\}_{i\in[n]}$ is much larger than differences among the squared norms, sketchy pivoting tends to behave like uniform sampling. Intuitively, there exists some $\left\{Z_{i}\right\}_{i\in[n]}$ with appropriate variance interpolating the two extreme cases such that sketchy pivoting mimics the behavior of squared-norm sampling and enjoys a nearly optimal skeleton complexity guarantee similar to Proposition 3.3.
In practice, sketchy pivoting (Algorithm 3) with Gaussian embedding provides high-quality skeleton selection with comparable error $\mathcal{E}_{\mathbf{X}}\left(S\right)$ to random pivoting (Algorithm 1, cf. Figure 1), but much more efficiently thanks to the parallelizability of sketching [16]. However, as a major empirical limitation, sketchy pivoting is not error-revealing and requires prior knowledge of the target rank $k$.
@@ 4 Robust Blockwise Random Pivoting
In light of the existing skeleton selection methods summarized in Table 1 that leverage/combine randomness and adaptiveness, a parallelizable, error-revealing, and exactly ID-revealing skeleton selection algorithm that attains similar skeleton and asymptotic complexities as the sequential random pivoting is a desirable missing piece.
Blockwise random pivoting is a natural extension of its sequential variation Algorithm 1 that consists of matrix-matrix multiplications and is therefore parallelizable. In the kernel formulation (with an input kernel matrix $\mathbf{X}\mathbf{X}^{\top}$), [5, Algorithm 2.2] introduced a blocked version of RPCholesky that can be generalized as blockwise random pivoting (BRP—Algorithm 4 with $\tau_{b}=0$). However, for a large block size $b\in\mathbb{N}$ that improves parallelizability, such plain BRP can pick up to $b$ times more skeletons than necessary due to the similar pitfall of squared-norm sampling instantiated in Example 2.4 (e.g. Figure 1). Moreover, in practice, the numerical instability issue tends to be exacerbated as the block size increases.
@@
4:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geq 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
@@
10:     $b^{\prime}\leftarrow\max\left\{i\in[b]\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geq\tau_{b}\cdot\left\|\mathbf{R}_{\mathbf{V}}\right\|_{F}^{2}\right\}$triangleright $O\left(b^{2}\right)$
@@
To resolve the challenges illustrated by Example 4.1, we introduce a robust blockwise random pivoting (RBRP) algorithm—Algorithm 4—that empirically achieves comparable skeleton complexities to the sequential adaptive methods (SRP and CPQR), as demonstrated in Figure 1 (right), while exploiting the parallelizability of matrix-matrix multiplications.
@@
In general, Algorithm 4 takes $O\left(ndk+bdk^{2}\right)$ time asymptotically (with a slightly larger lower-order term, cf. $O\left(ndk+dk^{2}\right)$ for the sequential version in Algorithm 1), where the dominant cost $O(ndk)$ consists of $k$ parallelizable passes through $\mathbf{X}$. Analogous to Algorithm 1, the storage of $\mathbf{L}^{(t)}$ and $\mathbf{Q}^{(t)}$ requires $O(nk)$ and $O(dk)$ memory, respectively. Meanwhile, the adaptiveness grants Algorithm 4 the blockwise error-revealing ability.
@@
CPQR Sequential greedy pivoting [22, Section 5.4.2]
SqNorm Squared-norm sampling Proposition 2.3 [13]
DPP DPP sampling [2, Equation 8][31, 21]
@@
SkCPQR Sketchy pivoting with CPQR Algorithm 3 [48]
SkLUPP Sketchy pivoting with LUPP Algorithm 3 [16]
@@
Figure 1: The skeletonization error $\mathcal{E}_{\mathbf{X}}\left(S\right)$ of algorithms in Table 2 on a GMM-based adversarial input as described in Example 4.1, where the adaptiveness is critical for achieving a good skeleton complexity/accuracy. (Left): Sequential random pivoting (SRP) and greedy pivoting (CPQR) enjoy the best skeleton complexities. The plain blockwise random pivoting (BRP) and greedy pivoting (BGP) tend to suffer from observably higher skeleton complexities than the sequential adaptive methods, while greedy selection tends to severely exacerbate such inefficiency. The squared-norm sampling (SqNorm) can be viewed as plain BRP with block size $b=k$, which suffers from a similar inefficiency as BRP/BGP. DPP sampling attains relatively good accuracy even without adaptiveness, thanks to its nearly optimal skeleton complexity [2, 25, 5], but with a much higher runtime in practice as a trade-off (e.g., with [21], k-DPP takes about $10^{3}\times$ longer than other algorithms in Table 2). (Right): With robust blockwise filtering (Remark 4.2), robust blockwise random pivoting (RBRP) and greedy pivoting (RBGP), as well as the sketchy pivoting algorithms, achieves the comparable, nearly optimal skeleton complexities as the sequential adaptive methods. It is worth highlighting that despite the comparable skeleton complexities between RBRP and RBGP, RBGP is generally much slower than RBRP (cf. Figure 2). This is because, without randomness, blockwise greedy pivoting tends to pick more redundant points in each block, which are later filtered out.
By comparing Figure 1 (left) and (right), we observe that the robust blockwise filtering (Remark 4.2) improves both blockwise random and greedy pivoting. In particular, with $\tau_{b}=\frac{1}{b}$, RBRP (and RBGP) empirically attains comparable skeleton complexities to those of the sequential random pivoting on the GMM adversarial input (Example 4.1). The same observation extends to all the natural and synthetic data matrices tested in our experiments (Section 6), which leads to the following 4.4:
@@
RBRP (Algorithm 4) with $\tau_{b}=\frac{1}{b}$ shares a skeleton complexity similar to sequential random pivoting (Proposition 3.3) in expectation.
@@
Consider Algorithm 4 at the $t$-th step with $\left|S\right|$ selected skeletons from the previous steps, $\left|S_{t}\right|=b$ candidates drawn from $\frac{\mathbf{d}^{(t-1)}}{\left\|\mathbf{d}^{(t-1)}\right\|_{1}}$, and $\left|S^{\prime}_{t}\right|=b^{\prime}$ remaining skeletons in $S_{t}$ that pass through the robust blockwise filtering (Remark 4.2). Now suppose we switch to sequential random pivoting after the $(t-1)$-th step and draw $b^{\prime}$ skeletons adaptively, denoted as $S^{\prime\prime}_{t}$. Then, the key rationale for 4.4 is that $\mathcal{E}_{\mathbf{X}}\left(S\cup S^{\prime}_{t}\right)\approx\mathcal{E}_{\mathbf{X}}\left(S\cup S^{\prime\prime}_{t}\right)$ with (a reasonably) high probability when $\tau_{b}=\frac{1}{b}$.
To see this, we observe the close connection between the selection scheme of $S^{\prime}_{t}$ and that of $S^{\prime\prime}_{t}$:
  • The robust blockwise filtering for $S^{\prime}_{t}$ in RBRP adaptively selects the point with the maximum residual norm in each step (i.e. CPQR on the residual matrix $\mathbf{V}$ in Algorithm 4), within a small subset $S_{t}$ of $b$ points sampled according to the squared-norm distribution over $[n]\setminus S$.
@@
Intuitively, every selection made for $S^{\prime}_{t}$ and $S^{\prime\prime}_{t}$ yields similar decay in the skeletonization error if the squared norms of the residual matrix, corresponding to remaining points in $S_{t}$, are around the maxima in the complement of the skeleton subset, so that these remaining points in $S_{t}$ can be selected by SRP with high probability666 It is worth mentioning that, in contrast to SRP, RBRP does not recompute the entire residual matrix after each skeleton selection. Instead, within each block, RBRP computes the residual matrix locally for $\mathbf{X}\left(S_{t},:\right)$ only, while updating the entire residual matrix at the end of each block via matrix-matrix multiplication. . Instead of computing the squared norms of the entire residual matrix explicitly for each selection as SRP (which leads to the undesired sequential updates), RBRP leverages the threshold $\tau_{b}=\frac{1}{b}$ to encourage the first $b^{\prime}$ points selected by the robust blockwise filtering in each block to have top squared norms in the corresponding residual matrices. Notice that with $\tau_{b}=\frac{1}{b}$, we enforce the sum of squared norms of the $b^{\prime}$ points to be larger than the average squared norm of the original block: $\left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geq\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$. Since the squared norms of residuals are non-increasing under QR updates, $\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$ serves as an overestimate for the squared norms of redundant points sampled in $S_{t}$.
As toy examples, if $\mathbf{V}$ is rank deficient (i.e. $\mathop{\mathrm{rank}}(\mathbf{V})=b^{\prime\prime}<b$), then there exists $b^{\prime}\leq b^{\prime\prime}$ such that $\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}:b,b^{\prime}:b\right)\right\|_{F}^{2}\geq\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$, but
$0=\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime\prime}+1:b,b^{\prime\prime}+1:b\right)\right\|_{F}^{2}\leq\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}+1:b,b^{\prime}+1:b\right)\right\|_{F}^{2}<\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}.$
Whereas if $\mathbf{V}\in\mathbb{R}^{d\times b}$ contains orthogonal columns with the same squared norm, then for all $i\in[b]$, we have $\left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geq\left|\mathbf{R}_{\mathbf{V}}\left(b,b\right)\right|^{2}=\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$, and therefore $b^{\prime}=b$.

### v5

@@ 3.1 Random Pivoting
For any $\epsilon>0$ and $r\in[n]$, the skeleton subset $S$ selected by Algorithm 1 provides $\mathbb{E}\left[\mathcal{E}_{\mathbf{X}}\left(S\right)\right]\leqslant\left(1+\epsilon\right)\left\|\mathbf{X}-{\langle\mathbf{X}\rangle}_{r}\right\|_{F}^{2}$ when333 Without strictly specifying, we generally assume that $\epsilon>0$ is small enough such that the $(r,\epsilon)$-ID is a non-trivial approximation, i.e. $\epsilon<\min\left\{1/\eta_{r},2^{r}\right\}$ such that both logarithmic terms are positive. Otherwise, the skeleton complexity guarantee still holds by replacing any negative logarithmic terms with zeros.
$k=\left|S\right|\geqslant\frac{r}{\epsilon}+r\cdot\min\left\{\log\left(\frac{1}{\epsilon\eta_{r}}\right),1+\log\left(\frac{2^{r}}{\epsilon}\right)\right\}.$
@@
Although random pivoting combines the strength of both adaptiveness and randomness and achieves an attractive skeleton complexity, the lack of hardware efficiency due to its sequential nature becomes a critical concern when considering empirical runtime.
@@
Given a sequence of matrix operations, it is well-known that an appropriate implementation using Level-3 BLAS, or matrix-matrix, operations will run more efficiently on modern processors than an optimal implementation using Level-2 or Level-1 BLAS [7]. This is largely due to the greater potential for the Level-3 BLAS to make more efficient use of memory caching in the processor. Notice that the computational bottleneck of Algorithm 1 with asymptotic complexity $O\left(ndk\right)$ consists of $k$ matrix-vector multiplications (Level-2 BLAS) with $\mathbf{X}$. This leads to the considerable slowdown of Algorithm 1 in practice.
@@ 3.2 Sketchy Pivoting
Alternative to random pivoting, randomness can be combined with adaptiveness through randomized linear embedding (also known as sketching), which leads to the sketchy pivoting methods [50, 62, 23, 28]. The general framework of sketchy pivoting [23, Algorithm 1] consists of two stages. In the first stage, randomness is incorporated through sketching, which serves two purposes.
  • From the computational efficiency perspective, sketching reduces the data dimension444 For applications like low-rank approximations and skeleton selection, constant oversampling like $l\geqslant k+10$ is usually sufficient in practice. We refer the interested readers to [37, 64, 52] for a general review of sketching. from $d$ to $l=O(k)$ and transfers the $O(ndk)$ computational bottleneck from sequential greedy pivoting to input-sparsity-time matrix-matrix multiplications.
@@
In the second stage, adaptiveness is utilized as greedy pivoting on the sketched sample matrix $\mathbf{Y}=\mathbf{X}\bm{\Omega}\in\mathbb{R}^{n\times l}$ to select skeletons. Both LUPP [33, Section 3.4] and CPQR [33, Section 5.4.2] on $\mathbf{Y}$ cost $O(nk^{2})$ asymptotically, whereas LUPP is remarkably faster in practice due to its superior parallelizability [23]. Intuitively, applying greedy pivoting on the top of sketching can be viewed as an analog of the squared-norm sampling with a slightly different squared-norm-based distribution.
Remark 3.3 (Sketchy pivoting v.s. random pivoting).
For example, consider a Gaussian random matrix $\bm{\Omega}$ with i.i.d. entries from $\mathcal{N}\left(0,1/l\right)$. Each row $\mathbf{y}_{i}=\bm{\Omega}^{\top}\mathbf{x}_{i}$ of $\mathbf{Y}$ is a Gaussian random vector with i.i.d. entries from $\mathcal{N}(\textbf{0},\left\|\mathbf{x}_{i}\right\|_{2}^{2}\mathbf{I}_{l})$. When applying greedy squared-norm pivoting (CPQR) in the second stage, the first pivot is selected according to $s_{1}\leftarrow\mathop{\mathrm{argmax}}_{i\in[n]}\left\|\mathbf{x}_{i}\right\|_{2}^{2}Z_{i}$ where every $Z_{i}=\sum_{j=1}^{l}Y_{ij}^{2}$ is a $\chi^{2}_{l}$ random variable but $\left\{Z_{i}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ i\in[n]\right\}$ are dependent. The intuition behind sketchy pivoting is similar to random pivoting, but the randomness arises from the random variables $Z_{i}$, instead of sampling.
In the extreme scenarios, when the variance of $\left\{Z_{i}\right\}_{i\in[n]}$ is negligible in comparison to the squared norms $\{\left\|\mathbf{x}_{i}\right\|_{2}^{2}\}_{i\in[n]}$, sketchy pivoting behaves similarly to greedy pivoting. Meanwhile, when the variance of $\left\{Z_{i}\right\}_{i\in[n]}$ is much larger than differences among the squared norms, sketchy pivoting tends to behave like uniform sampling. Intuitively, there exists some $\left\{Z_{i}\right\}_{i\in[n]}$ with appropriate variance interpolating the two extreme cases such that sketchy pivoting mimics the behavior of squared-norm sampling and enjoys a nearly optimal skeleton complexity guarantee similar to Proposition 3.1.
In practice, sketchy pivoting provides high-quality skeleton selection with comparable error $\mathcal{E}_{\mathbf{X}}\left(S\right)$ to random pivoting (Algorithm 1, cf. Figure 1), but much more efficiently thanks to the parallelizability of sketching [23]. However, as a major empirical limitation, sketchy pivoting is not rank-adaptive and requires prior knowledge of the target rank $k$.
@@ 4 Robust Blockwise Random Pivoting
In light of the existing skeleton selection methods summarized in Table 1, a hardware-efficient, rank-adaptive, and ID-revealing skeleton selection algorithm that attains similar skeleton and asymptotic complexities as the sequential random pivoting raises as a desirable missing piece.
Blockwise random pivoting is a natural extension of its sequential variation Algorithm 1 that consists of matrix-matrix multiplications and is, therefore, hardware-efficient. In the kernel formulation (with an input kernel matrix $\mathbf{X}\mathbf{X}^{\top}$), [8, Algorithm 3] introduced a blocked version of RPCholesky that can be generalized as blockwise random pivoting (BRP—Algorithm 2 with $\tau_{b}=0$). However, when the block size $b\in\mathbf{N}$ is taken to be large enough to fully realize the advantages of hardware efficiency, such plain BRP can have up to a $b$ times larger skeleton complexity than necessary due to the similar pitfall of squared-norm sampling instantiated in Example 2.3 (e.g. Figure 1).
@@
4:$\mathbf{d}^{(0)}\in\mathbb{R}_{\geqslant 0}^{n}$ with $\mathbf{d}^{(0)}(i)=\left\|\mathbf{x}_{i}\right\|_{2}^{2}\leavevmode\nobreak\ \forall\leavevmode\nobreak\ i\in[n]$triangleright $O(nd)$
@@
10:     $b^{\prime}\leftarrow\max\left\{i\in[b]\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geqslant\tau_{b}\cdot\left\|\mathbf{R}_{\mathbf{V}}\right\|_{F}^{2}\right\}$triangleright $O\left(b^{2}\right)$
@@
To resolve the challenges illustrated by Example 4.1, we introduce a robust blockwise random pivoting (RBRP) algorithm—Algorithm 2—that empirically achieves comparable skeleton complexities to the sequential adaptive methods (SRP and CPQR), as demonstrated in Figure 1 (right), while exploiting the hardware efficiency of matrix-matrix multiplications.
@@
The worst-case asymptotic complexity of Algorithm 2 is $O\left(ndk+bdk^{2}\right)$ (with a slightly larger lower-order term, cf. $O\left(ndk+dk^{2}\right)$ for SRP in Algorithm 1) when each step of robust blockwise filtering in Algorithm 2 only keeps a single skeleton. In practice, robust blockwise filtering typically retains a reasonable portion of skeletons in each block, leading to a time complexity close to $O\left(ndk+dk^{2}\right)$ (see Section 6.2). The dominant cost $O(ndk)$ consists of $k$ hardware-efficient passes through $\mathbf{X}$ as matrix-matrix multiplications. Analogous to Algorithm 1, the storage of $\mathbf{L}^{(t)}$ and $\mathbf{Q}^{(t)}$ requires $O(nk)$ and $O(dk)$ memory. The adaptive nature of Algorithm 2 further facilitates the rank-adaptiveness.
@@
CPQR Sequential greedy pivoting [33, Section 5.4.2]
SqNorm Squared-norm sampling (5[20]
DPP DPP sampling [5, Equation 8][45, 31]
@@
SkCPQR Sketchy pivoting with CPQR [62, Algorithm 4]
SkLUPP Sketchy pivoting with LUPP [23, Algorithm 1]
@@
Figure 1: The relative skeletonization error $\mathcal{E}_{\mathbf{X}}\left(S\right)/\|\mathbf{X}\|_{F}^{2}$ and skeleton selection time (excluding interpolation matrix construction) of algorithms in Table 2 on the adversarial GMM matrix described in Section 6.1 constructed according to Example 4.1. For the sketchy pivoting methods, we only show the results of SkLUPP, as SkCPQR is known to have similar skeleton complexities as SkLUPP but with higher runtimes [23].
Figure 1 concretizes the adversarial input in Example 4.1 and demonstrates the empirical effectiveness of RBRP for skeleton selection.
  • In particular, the skeletonization errors on the left unveil the failure of plain BRP/BGP in terms of large skeleton complexities due to the lack of adaptiveness. Moreover, greedy selection in BGP tends to severely exacerbate this problem. SqNorm can be viewed as plain BRP with block size $b=k$ and therefore suffers from similar large skeleton complexities as BRP/BGP, despite its low asymptotic complexity of $O(nd)$.
  • DPP sampling attains relatively good accuracy even without adaptiveness, thanks to its nearly optimal skeleton complexity [5, 36, 8], but with much higher runtimes as a trade-off (e.g., with [31], k-DPP takes about $10^{3}\times$ longer than other algorithms in Table 2).
  • Providing even better skeleton complexities than DPP, SRP and CPQR achieve one of the best skeletonization errors thanks to their adaptiveness. However, the sequential nature of SRP and CPQR leads to higher runtimes than the blockwise methods.
  • With robust blockwise filtering (Remark 4.2), RBRP achieves comparable skeleton complexities to SRP and CPQR, while maintaining the hardware efficiency of blockwise methods. Notice from the runtimes on the right that despite the comparable skeleton complexities to RBRP, RBGP is slower than RBRP on the adversarial GMM matrix. This is because RBGP with greedy selection tends to pick more redundant points in each block, which are later removed by the robust blockwise filtering.
  • For skeleton selection only, SkLUPP provides the state-of-the-art performance — one of the best skeletonization errors with the most competitive runtimes. However, the lack of rank-adaptiveness and ID-revealing property becomes a major drawback for SkLUPP as an ID algorithm (see Section 5.2).
Beyond the GMM adversarial input, extensive numerical experiments in Section 6 show that RBRP with $\tau_{b}=\frac{1}{b}$ empirically attains comparable skeleton complexities as sequential random pivoting, which leads to the following conjecture.
@@
RBRP (Algorithm 2) with $\tau_{b}=\frac{1}{b}$ has similar skeleton complexity as sequential random pivoting (Proposition 3.1) in expectation.
@@
Consider Algorithm 2 at the $t$-th step with $\left|S\right|$ selected skeletons from the previous steps, $\left|S_{t}\right|=b$ candidates drawn from ${\mathbf{d}^{(t-1)}}/{\left\|\mathbf{d}^{(t-1)}\right\|_{1}}$, and $\left|S^{\prime}_{t}\right|=b^{\prime}$ remaining skeletons in $S_{t}$ that pass through the robust blockwise filtering (Remark 4.2). Now suppose we switch to sequential random pivoting after the $(t-1)$-th step and draw $b^{\prime}$ skeletons adaptively, denoted as $S^{\prime\prime}_{t}$. Then, the key rationale for 4.4 is that $\mathcal{E}_{\mathbf{X}}\left(S\cup S^{\prime}_{t}\right)\approx\mathcal{E}_{\mathbf{X}}\left(S\cup S^{\prime\prime}_{t}\right)$ with (a reasonably) high probability when $\tau_{b}=\frac{1}{b}$.
On this, we observe the close connection between the selection scheme of $S^{\prime}_{t}$ and that of $S^{\prime\prime}_{t}$:
  • The robust blockwise filtering for $S^{\prime}_{t}$ in RBRP adaptively selects the point with the maximum residual norm in each step (via CPQR on $\mathbf{V}$ in Algorithm 2), within a small subset $S_{t}$ of $b$ points sampled according to the squared-norm distribution over $[n]\setminus S$.
@@
Intuitively, every selection made for $S^{\prime}_{t}$ and $S^{\prime\prime}_{t}$ yields similar decay in the skeletonization error if the squared norms of the residual matrix, corresponding to remaining points in $S_{t}$, are around the maxima in the complement of the skeleton subset so that these remaining points in $S_{t}$ can be selected by SRP with high probability555 It is worth mentioning that, in contrast to SRP, RBRP does not recompute the entire residual matrix after each skeleton selection. Instead, within each block, RBRP computes the residual matrix locally for $\mathbf{X}\left(S_{t},:\right)$ only, while updating the entire residual matrix at the end of each block via matrix-matrix multiplication. . Instead of computing the squared norms of the entire residual matrix explicitly for each selection as SRP (which leads to the undesired sequential updates), RBRP leverages the threshold $\tau_{b}=\frac{1}{b}$ to encourage the first $b^{\prime}$ points selected by the robust blockwise filtering in each block to have top squared norms in the corresponding residual matrices. Notice that with $\tau_{b}=\frac{1}{b}$, we enforce the sum of squared norms of the $b^{\prime}$ points to be larger than the average squared norm of the original block: $\left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geqslant\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$. Since the squared norms of residuals are non-increasing under QR updates, $\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$ serves as an overestimate for the squared norms of redundant points sampled in $S_{t}$.
As toy examples, if $\mathbf{V}$ is rank deficient (i.e. $\mathop{\mathrm{rank}}(\mathbf{V})=b^{\prime\prime}<b$), then there exists $b^{\prime}\leqslant b^{\prime\prime}$ such that $\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}:b,b^{\prime}:b\right)\right\|_{F}^{2}\geqslant\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$, but
$0=\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime\prime}+1:b,b^{\prime\prime}+1:b\right)\right\|_{F}^{2}\leqslant\left\|\mathbf{R}_{\mathbf{V}}\left(b^{\prime}+1:b,b^{\prime}+1:b\right)\right\|_{F}^{2}<\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}.$
Whereas if $\mathbf{V}\in\mathbb{R}^{d\times b}$ contains orthogonal columns with the same squared norm, then for all $i\in[b]$, we have $\left\|\mathbf{R}_{\mathbf{V}}\left(i:b,i:b\right)\right\|_{F}^{2}\geqslant\left|\mathbf{R}_{\mathbf{V}}\left(b,b\right)\right|^{2}=\frac{1}{b}\left\|\mathbf{V}\right\|_{F}^{2}$, and therefore $b^{\prime}=b$.